OpenStudy (anonymous):

Which of the following ordered pairs is not a solution to the inequality y <= -2x - 3? (-4, 0) (-1, 0) (-2, -1) (0, -4)

OpenStudy (anonymous):

Plug in the x and y values and see if it satisfies the inequality.. So for the fist point (-4,0) \[0 \le -2(-4) - 3\]\[0 \le 8 - 5\]

OpenStudy (anonymous):

\[0 \le 3\] This is true so (-4,0) is a solution

OpenStudy (anonymous):

So we have to move to the next pair \[0 \le -2(-1) - 3\] \[0 \le 2 - 3\] \[0 \le -1\] This is not true therefore (-1,0) is not a solution.